OpenStudy (lgbasallote):

how many d orbitals have n = 6? can someone tell me how to get this?

OpenStudy (chmvijay):

20 d orbitals

OpenStudy (lgbasallote):

*how*

OpenStudy (lgbasallote):

and btw 20 is wrong

OpenStudy (chmvijay):

just see in this image for n =6 u get 4 d orbits 3d,4d,5d and 6d, each d orbit contains 5d orbitals hence 20 d orbitals u get

OpenStudy (chmvijay):

if u want interms of orbits it will be four

OpenStudy (lgbasallote):

it's actually 5..

OpenStudy (chmvijay):

how tell me

OpenStudy (anonymous):

Remember that 4-part quantum descriptor I posted yesterday for you? ;-) \[\large( n , \ell , m_\ell , m_s )\] Well... \(\huge\cdot\) The value for \(\large \ell\) determines an orbital's shape \(\huge\cdot\) The value for \(\large m_\ell\) its orientation; but because some orbitals are described by complex numbers (that a+b\(i\) stuff) it sometimes has an effect on shape too How to get this? Um, I'm going to go ahead an assume you're talking in generality because the "how" is a large topic of a whole science called quantum mechanics. I am not going to give you a entire written description of how quantum mechanics works because of some very good reasons: 1. I don't have the time, and the rewards for all that effort doesn't justify it. I'm willing to help, but you need to narrow the focus of your question, like a lot. Ask something specific 2. I don't personally feel qualified to even consider myself an expert on quantum mechanics. And quantum mechanics itself is incredibly complex, being research by people farm more experienced and intelligent than myself. 3. Quantum mechanics deals with stuff that can't be seen directly and is mostly theoretical (i.e.: untested or unobserved in real-world experiment) or that experimental data is based on the best estimations we can do at this time, (i.e.: compare \(\pi\cdot \sqrt{-2}\) to a decimal? haha ok best of luck to you with that one). We're starting to find that the universe is built on special types of complex numbers and that our "convenient" decimal number system and Cartesian coordinates are an oversimplified version of the TRUE complexity of the universe. 4. I would probably encounter the character limit. In my response. ^ read me please. With that said, D-orbitals are a combination of the following based on electron pairs (which have opposite spins). What does this look like visually? Glad you asked, and thank you Wikimedia for having this: http://upload.wikimedia.org/wikipedia/commons/thumb/e/e1/D_orbitals.svg/2000px-D_orbitals.svg.png The also have a computer approximation of what n={1 to 6} for hydrogen (1s\(^1\)) where you'd be most likely to find the electron: http://en.wikipedia.org/wiki/File:HydrogenOrbitalsN6L0M0.png Keep in mind electrons sometimes are like Cartman from Southpark going, "F U guys I'm going home!" (this is formally known has the Heisenberg Uncertainty Principle) It's a probability the electron will be there, not a guarantee. So... If you look at the image, it's 4+4+4+4+4 = 4 * 5 = 20 orbital "lobes"

OpenStudy (lgbasallote):

it's still a while ago for me haha

OpenStudy (anonymous):

Again @lgbasallote , you'd do well to narrow the focus of your question ;-)

OpenStudy (anonymous):

Makes me wonder if you actually read what I answered for you yesterday too :-(

OpenStudy (lgbasallote):

well i just wrote what's in the paper...that's the exact question and yes i read it...i got some of it...some didnt register..sadly

OpenStudy (chmvijay):

haaaa nice!!!

OpenStudy (lgbasallote):

exact question is how many d orbitals have n = 6? a) 2 b) 5 c) 10 d) 7 e) 18

OpenStudy (lgbasallote):

and i also dont get what quantum mechanics has to do with this..

OpenStudy (anonymous):

Ah you're probably looking at 6d\(^{something}\)

OpenStudy (lgbasallote):

could be..

OpenStudy (chmvijay):

it hink its only 6d then i can say any Nd orbit will have 5 orbitals!!!!

OpenStudy (lgbasallote):

wait no...the answer is 5

OpenStudy (lgbasallote):

how did you get it @chmvijay ?

OpenStudy (anonymous):

See? The way the question is worded, and posting the choice option if it's multiple choice is VERY necessary on a site like this. We don't have the problem. We have to guess the subject matter when thinking of an answer. For this I would refer you to the periodic table... (hold on I'll get a good link)

OpenStudy (chmvijay):

hey it is not for 6 d if you consider any d orbit it contains definately 5 orbitals

OpenStudy (lgbasallote):

oh btw.. @agentx5 no offense..you have a habit of using complex words :S im not experienced in chemistry so i dont know your terms sometimes

OpenStudy (chmvijay):

question setting has little bit not understandable , since 2o is not there we should go for 5 i think their intension of asking might be different

OpenStudy (lgbasallote):

i didnt know choices were important here lol

OpenStudy (anonymous):

http://www.meta-synthesis.com/webbook/34_qn/janet_number.png D = 5 orbital combinations, 2 pairs of electrons per orbital, in the 5 formations I showed you above in my first mega-post For example can you name how many orbital combinations you'd find for the F-orbital block?

OpenStudy (chmvijay):

seven

OpenStudy (lgbasallote):

oh wait does it come from -2, -1, 0, 1, 2? that thingy?

OpenStudy (chmvijay):

no no i dont think so

OpenStudy (anonymous):

@chmvijay is correct! You have 7 different formations for orbitals in the F-block ;-) 14 columns / 2 pairs of electrons = 7 formations

OpenStudy (anonymous):

Isn't the periodic table handy? ;-)

OpenStudy (lgbasallote):

but it's d block l

OpenStudy (lgbasallote):

lol*

OpenStudy (anonymous):

Yes and if you're filling electrons in the D-block, how many combinations are there by the time you've got to 3d\(^{10}\) (for understanding, can you figure out which element that is, if you're assuming it's neutral/uncharged?)

OpenStudy (lgbasallote):

i dont know what you're talking about sorry

OpenStudy (lgbasallote):

am i right though? the answer is 5 because there are 5 orientations for d?

OpenStudy (lgbasallote):

i have to go sleep now =_= ill come back to this in the morning....

OpenStudy (anonymous):

For D orbitals: 5 orbital orientations (again, the wikipedia picture diagram) 10 electrons to fill a D-orbital up completely (5 * 2 e- per pair) 20 orbital lobes (quantum probability curves)

OpenStudy (anonymous):

I agree 5 is most likely the best answer for how the question was worded. Try it :-)

OpenStudy (anonymous):

A "d" orbital has angular momentum quantum number l = 2. (s = 0, p = 1, d = 2, f = 3, g = 4 and so forth). The number of d orbitals is given by how many values of the azimuthal quantum number m_l there are. m_l always goes from -l to +l, e.g. in the case of l = 2 from -2 to +2 (-2,-1,0,+1,+2), for a total of 2*l + 1, e.g. in the case of l=2 a total of 5 orbitals. The value of the principal quantum number n doesn't matter except that it must be greater than 2, because n = 1 and n = 2 don't have d orbitals at all. (The allowed values of l are 0 to n-1.)

OpenStudy (anonymous):

@Carl_Pham I got a syntax helpful tip for ya! :-D \ell = \(\ell\) subscript\large_{\ell} = subscript\(\large_{\ell}\) He_2^4 = He\(_2^4\) To make the syntax active code you need to put a \( before it

OpenStudy (anonymous):

and a \) after it (I can't put this in the same post or it activates)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Give it a try @Carl_Pham :-D

OpenStudy (anonymous):

It's incredible helpful for writing stuff in chemistry and physics, two sections you seem to frequent.

OpenStudy (lgbasallote):

okay thanks..it's beginning to make some sense now :) just a few more practice!! \m/